OpenStudy (anonymous):

Solve by elmination/ substitution 4x-2y+5z=6 3x+3y+8z=4 x-5y-3z=5

OpenStudy (anonymous):

Do you solve for x using the third equation?

OpenStudy (anonymous):

you can then put that answer for x into both of the first two equations and then solve for y or z then put your answer back into the original equations

OpenStudy (anonymous):

this works out good so you get x=5y+3z+5 so now sub for x in other two equations and substract those two equations and the y 's cancell and that gives you z and the go backwards.

OpenStudy (anonymous):

Wouldn't the z cross out too? 4(5y+3z=5)-2y+5z=6 20y+12z+20-2y+5z+6 18y+17z+20=6 18y+17z=-14 3(5y+3z+5)+3y+8z=4 15y+9z+15+3y+8z=4 18y+17z+15=4 18y+17z=-11

OpenStudy (anonymous):

yes let me look at something

OpenStudy (anonymous):

so there is no solutions

OpenStudy (anonymous):

Thank you very much. :)